Articles

1. CHARM II Collaboration, P. Vilain et al. Neutral current coupling constant from neutrino- and antineutrino-electron scattering, Phys. Lett. B281 (1992) 159-166.

2. CHARM II Collaboration, D. Geiregat et al. Calibration and performance of the CHARM II detector, Nucl. Instr. & Meth. A325 (1993) 92-108.

3. CHARM II Collaboration, P. Vilain et al. Measurement of differential cross sections for muon-neutrino electron scattering, Phys. Lett. B302 (1993) 351-355.

4. CHARM II Collaboration, M. Gruwé et al. Search for ν_{μ} → ν_{τ} oscillation, Phys. Lett. B309 (1993) 463-468.

5. CHARM II Collaboration, P. Vilain et al. Coherent single charged pion production by neutrinos, Phys. Lett. B313 (1993) 267-275.

6. CHARM II Collaboration, P. Vilain et al. Flavour Universality of Neutrino Couplings with the Z, Phys. Lett. B320 (1994) 203-205.

7. S. Buontempo et al.

Construction and test of calorimeter modules for the CHORUS experiment, Nucl. Instr. & Meth. A349 (1994) 70-80.

8. CHARM II Collaboration, P. Vilain et al. Constraints on additional Z bosons from neutrino electron scattering measurements, Phys. Lett. B332 (1994) 465-470.

9. CHARM II Collaboration, P. Vilain et al. Precision Measurement of Electroweak Parameters from the Scattering of Muon-Neutrinos on Electrons, Phys. Lett. B335 (1994) 246-252.

10. CHARM II Collaboration, P. Vilain et al. Search for muon to electron neutrino oscillations, Z. Phys. C64 (1994) 539-544.

11. CHARM II Collaboration, P. Vilain et al. Search for Heavy Isosinglet Neutrinos, Phys. Lett. B343 (1995) 453-458, Phys. Lett. B351 (1995) 387-392.

12. CHARM II Collaboration, P. Vilain et al. Experimental study of the electromagnetic properties of the muon neutrino in neutrinoelectron scattering, Phys. Lett. B345 (1995) 115-118.

13. A. Artamonov et al.

A liquid scintillator calorimeter for the forward region of a LHC experiment, Nucl. Instr. & Meth. A362 (1995) 386-394.

14. CHARM II Collaboration, P. Vilain et al. A precise measurement of the cross section of the inverse muon decay ν_{μ} + e^{−} → μ^{−} + ν_{e}, Phys. Lett. B364 (1995) 121-126.

15. E. Di Capua et al.

Response to electrons and pions of the calorimeter for the CHORUS experiment, Nucl. Instr. & Meth. A378 (1996) 221-232.

16. CHORUS Collaboration, E. Eskut et al. The CHORUS experiment to search for ν_{μ} → ν_{τ} oscillation, Nucl. Instr. & Meth. A401 (1997) 7-44.

17. CHORUS Collaboration, E. Eskut et al. A Search for ν_{μ} → ν_{τ} Oscillation, Phys. Lett. B424 (1998) 202.

18. CHARM II Collaboration, P. Vilain et al. Experimental search for muonic photons, Phys. Lett. B434 (1998) 200.

19. CHORUS Collaboration, E. Eskut et al. A Search for ν_{μ} → ν_{τ} oscillation using the τ decay modes into a single charged particle, Phys. Lett. B434 (1998) 205.

20. CHORUS Collaboration, P. Annis et al. Observation of neutrino induced diffractive D_{s}^{∗+} production and subsequent decay D_{s}^{∗+} → D_{s}^{+} → τ^{+} → μ^{+} , Phys. Lett. B435 (1998) 458.

21. CHARM II Collaboration, P. Vilain et al. Leading-order QCD analysis of neutrino-induced dimuon events, CERN-EP/98-128, August 3, 1998. Submitted to Eur. Phis. J. C.

Proposals

1. CHORUS Collaboration, M. de Jong et al. A new search for ν_{μ} → ν_{τ} oscillation, CERN-PPE/93-131 July 20, 1993.

2. R. van Dantzig et al.

High resolution tracking devices based on capillaries filled with liquid scintillator, CERN/LHCC 95-7, P60/LDRB, February 27, 1994.

3. A. S. Ayan et al.

A high sensitivity short baseline experiment to search for ν_{μ} → ν_{τ} oscillation, CERN-SPSC/97-5, SPSC/I 213, March 14, 1997.

4. H. Shibuya et al.

The OPERA emulsion detector for a long baseline neutrino oscillation experiment, LNGS-LOI 8/97, CERN-SPSC 97-24, SPSC/I 218, October 1997.

5. A. S. Ayan et al.

TOP: A Prototype for the TOSCA Experiment, CERN-SPSC/98-02, January 1998.

6. K. Kodama et al.

The OPERA ν_{τ} appearance experiment in the CERN-Gran Sasso neutrino beam, LNGS-LOI 8/97-Add. 1, CERN-SPSC 98-25, SPSC/M612, October 9, 1998.

Conference contributions

1. E. Di Capua et al.

Design and test performance of the high resolution calorimeter of the CHORUS experiment, presented by G. Fiorillo at the 1992 Meeting of the Italian Physical Society, Pavia (Italy), October 5-10 1992. Published in the conference Proceedings.

2. CHARM II Collaboration, P. Vilain et al. Measurement of electroweak parameters from ν_{μ} − e^{−} scattering, presented by A. G. Cocco at the “XXVII Rencontre de Moriond”, Les Arcs (France), March 15-22 1992. Proceedings edited by J. Trân Thanh Vân, Ed. Frontières, Gif sur Yvette (1993) 81-87.

3. CHARM II Collaboration, P. Vilain et al. Electroweak paramenters from neutrino-electron scattering, presented by F. Marchetti-Stasi at the “4th S. Miniato Topical Seminar on The Standard Model and just bejond”, S. Miniato (Italy), June 1-5 1992. Proceedings edited by F. Navarria and P. G. Pelfer, World Scientific (1993).

4. S. Buontempo et al.

The high resolution calorimeter of the CHORUS experiment, presented by S. Buontempo at the “3rd International Conference on Advanced Technology and Particle Physics”, Como (Italy), June 22-26 1992. Published in Nucl. Phys. B (Proc. Suppl.) 32 (1993) 70-76.

5. S. Buontempo et al.

Design and test performance of the high resolution CHORUS calorimeter, presented by A. Ereditato at the “3rd International Conference on Calorimetry in High Energy Physics”, Chorpus Christi (Texas), September 29–October 2, 1992. Proceedings edited by P. Hale and J. Siegrist, World Scientific (1993).

6. S. Buontempo et al.

Construction and test of modules of the CHORUS calorimeter, presented by G. Fiorillo at the “5th S. Miniato Topical Seminar on Experimental Apparatus for High Energy Physics and Astrophysics”, S. Miniato (Italy), April 26-30 1993. Published in Nucl. Instr. & Meth. A344 (1994) 85-91.

7. E. Di Capua et al.

The lead/scintillating fiber and strip calorimeter of the CHORUS experiment, presented by A. Ereditato at “SCIFI ‘93”, Notre Dame (Indiana), October 24-28 1993. Proceedings edited by A.D. Bross, R.C. Ruchti, M.R. Wayne, World Sci. (1995) 215-229.

8. CHARM II Collaboration, P. Vilain et al. Tests of the Standard Model from ν_{μ} − e^{−} scattering, presented by A. G. Cocco at the “XXIX Rencontre de Moriond”, Meribel (France), March 12-19 1994. Proceedings edited by J. Trân Thanh Vân, Ed. Frontières, Gif sur Yvette (1994)M82 23-29.

9. A. Artamonov et al.

Liquid scintillator calorimetry for the LHC, presented by S. Buontempo at the “6th Pisa Meeting on Advanced Detectors”, La Biodola, Isola d’Elba (Italy), May 22-28 1994. Published in Nucl. Instr. & Meth. A360 (1995) 240-244.

10. S. Buontempo et al. The Chorus Calorimeter: Test Results, presented by G. Fiorillo at the “6th Pisa Meeting on Advanced Detectors”, La Biodola, Isola d’Elba (Italy), May 22-28 1994. Published in Nucl. Instr. & Meth. A360 (1995) 197-200.

11. CHARM II Collaboration, P. Vilain et al. Flavour Universality of Neutrino Couplings with the Z, presented by A. G. Cocco at the “XVII International Conference on High Energy Physics”, Glasgow (Scotland), July 20-27 1994. Proceedings edited by P. J. Bussey, I. G. Knowles, Bristol Institute of Physics, 1995.

12. A. Artamonov et al.

Test beam results from a very forward liquid scintillator calorimeter for the LHC, presented by F. Garufi at the “5th International Conference on Calorimetry in High Energy Physics”, Brookhaven (USA), September 1994. Proceedings edited by H. A. Gordon and D. Rueger. World Scientific (1995).

13. A. Artamonov et al.

A very forward calorimeter for the LHC: experimental results, presented by S. Buontempo at the “4th International Conference on Advanced Technology and Particle Physics”, Como (Italy), October 3-7 1994. Published in Nucl. Phys. B (Proc. Suppl.) 44 (1995) 40-44.

14. CHARM II Collaboration, P. Vilain et al. A precise measurement of the cross section of the inverse muon decay ν_{μ} + e^{−} → μ^{−} + ν_{e}, presented by G. Fiorillo at the “International Europhysics Conference on High Energy Physics”, Brussels (Belgium), July 27 - August 2 1995. Proceedings edited by j. lemonne, C. Vander Velte and F. Verbeure, World Sci. (1996) 518-519.

15. CHARM II Collaboration, P. Vilain et al. Neutrino–lepton scattering in the CHARM II experiment, presented by F. Marchetti-Stasi at the “3rd Tallin Symposium on Neutrino Physics, Lohusalu (Estonia), October 1995. Proceedings edited by I. Ots, J. Lõhmus, P. Helde, L. Palgi. Tartu 1995.

16. E. Di Capua et al.

Performance of the CHORUS lead/scintillating fiber calorimeter, presented by F. Marchetti-Stasi at the “6th S. Miniato Topical Seminar on Experimental Apparatus for High Energy Physics and Astrophysics”, S. Miniato (Italy), May 1996. To be published in the proceedings.

17. P. Annis et al.

Capillary detectors for high resolution tracking, presented by F. Garufi at the “6th S. Miniato Topical Seminar on Experimental Apparatus for High Energy Physics and Astrophysics”, S. Miniato (Italy), May 1996. To be published in the proceedings.

18. S. Buontempo et al.

The CHORUS calorimeter: test beam results and operation with neutrino beams, presented by P. Migliozzi at the “VI International Conference on Calorimetry in High-Energy Physics”, Frascati (Roma), 8-14 june 1996 Proceedings edited by A. Antonelli, S. Bianco, A. Calcaterra, F.L. Fabbri, Frascati Physics Series, Frascati (Italy), 1996.

19. CHORUS Collaboration, Results of the CHORUS neutrino oscillation experiment, presented by P. Migliozzi at the “XI Rencontres de Physique de la Vallee d’Aoste”, La Thuile (Aosta), 3-8 March 1997 To be publishedin the proceedings.

20. CHORUS Collaboration, E. Eskut et al. Results on ν_{μ} → ν_{τ} oscillations from the CHORUS experiment, presented by P. Migliozzi at the “XVIII International Symposium on Lepton Photon Interactions”, Hamburg, Germany, July 28 - August 1, 1997. To be published in the proceedings.

21. CHORUS Collaboration, E. Eskut et al. New results from the CHORUS neutrino oscillation experiment, presented by P. Migliozzi at the “XXIX International Conference on High Energy Physics”, Vancouver, Canada, July 23–29, 1998. To be published in the proceedings.

22. CHORUS and NOMAD Collaborations Recent results from neutrino oscillation experiments at CERN, presented by G. Fiorillo at the “Second Meeting on New Worlds in Astroparticle Physics”, Faro, Portugal, September 3–5, 1998. To be published in the proceedings.